OpenStudy (anonymous):

FAN and MEDAL 27x^z/-3x^2z^

OpenStudy (anonymous):

\[\frac{ 27x ^{2}z }{ -3x ^{2}z ^{4}}\]

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

There is a really helpful math site that will give you basically all the answers. its https://www.mathway.com/ the answer should be \[-9/z ^{3}\]

OpenStudy (anonymous):

To work with exponents, take the variable's larger exponent and subtract the smaller exponent for the corresponding variables: \[\frac{ 27x ^{2} z }{ -3x ^{2} z ^{4} }\]\[\frac{ 27x ^{2-2} }{-3 z ^{4-1} }\]\[\frac{ -9 }{ z ^{3}}\] For x^2, they canceled because they have the same value on top and bottom, 2-2=0; anything to the 0 power is 1, so remove x from the equation. For z, the z^4 on the bottom is bigger than the z on the numerator, so the z on top is raised to the 1st power, subtract that from the power of 4, and you get z^3 on the denominator.